# OT Score: An OT based Confidence Score for Unsupervised Domain Adaptation

### v1

@@ Appendix B Proofs
Now consider the optimal transport map $T_{\nu}^{\mu}$ restricted on $(E_{\nu_{1}}\cap F\cap B_{R})\cup(E_{\nu_{2}}\cap F)$. By [30, Theorem 4.6], this restricted map is an optimal transport map between the marginal measures.
@@
[…]
@@
Choosing the difference $C-D$ allows us to conclude $(T_{\bar{\nu}}^{\hat{\mu}})_{\#}\bar{\nu}_{1}=\hat{\mu}_{1}$ and $(T_{\bar{\nu}}^{\hat{\mu}})_{\#}\bar{\nu}_{2}=\hat{\mu}_{2}$ by setting $\bm{w}=[\bm{w}_{1},\bm{w}_{2}]$.
Conversely, let $\bm{w}$ be the corresponding weight vector of $T_{\bar{\nu}}^{\hat{\mu}}$ and assume $(T_{\bar{\nu}}^{\hat{\mu}})_{\#}\bar{\nu}_{1}=\hat{\mu}_{1}$, $(T_{\bar{\nu}}^{\hat{\mu}})_{\#}\bar{\nu}_{2}=\hat{\mu}_{2}$. Then $\bm{w}_{1}$ (or $\bm{w}_{2}$) differs from $\bm{m}$ (or $\bm{l}$) by some constant $C$ (or $D$) [8, Theorem 2]. By Theorem 6,
@@
Appendix C SDOTDA
C.1 Semi-discrete OT for Domain Adaptation
We treat the source distribution as a discrete probability measure and the target distribution as a continuous probability measure that is absolutely continuous with respect to the Lebesgue measure. SDOT-DA assigns labels to target samples based on the semi-discrete OT map from target features to source features, using the labels of their transport destinations.
Algorithm 2 SDOTDA
1:Input: Pre-trained model $f_{\theta}=h_{\bm{w}}\circ\phi_{\bm{v}}$, discrete probability measure $\hat{\mu}$ constructed from $N_{\mathcal{S}}$ source feature samples $\phi_{\bm{v}}(\textbf{X}_{\mathcal{S}})$, source data labels $\textbf{Y}_{\mathcal{S}}$, target data samples $\textbf{X}_{\mathcal{T}}$, entropic regularization parameter $\varepsilon$, learning rate $\gamma$.
2:Compute target features $\nu:=(\phi_{\bm{v}})_{\#}\nu_{\mathcal{T}}.$
3:for $iteration=1,2,\ldots,max\_iter$ do
4:     Draw a sample $x$ from $\nu$.
5:     For each sample $z_{j}\in\hat{\mu}$, compute the smoothed version of the indicator function of the Laguerre cell $\mathbb{L}_{\bm{w}}(z_{j})$: $\chi_{j}^{\varepsilon}(x,\bm{w})=\frac{e^{\frac{-\|x-z_{j}\|+\bm{w}_{j}}{\varepsilon}}}{\sum_{\ell}e^{\frac{-\|x-z_{\ell}\|+\bm{w}_{j}}{\varepsilon}}}$.
6:     Update $\bm{w}$ with the stochastic gradient at $x$: $\bm{w}=\bm{w}-\gamma\left(\chi_{j}^{\varepsilon}(x,\bm{w})-\frac{1}{N_{\mathcal{S}}}\right)_{j=1}^{N_{\mathcal{S}}}\in\mathbb{R}^{N_{\mathcal{S}}}$.
7:end for
8:for $x\in\textbf{X}_{\mathcal{T}}$ do
9:     Compute $j=\operatorname*{arg\,max}_{z_{j}\in\hat{\mu}}\chi_{j}^{\varepsilon}(x,\bm{w})$
10:     Label of $x$leftarrow label of $z_{j}$
11:end for
Remark 5.
An advantage of SDOTDA is that once the weight vector $\bm{w}$ is computed, we can easily classify new data points from the target distribution using the (approximated) indicator function without recomputing the semi-discrete OT. Also, this algorithm does not require to re-train the neural network so it is preferred when the target domain dataset is not sufficient for training a neural network.
As mentioned in section 4.2, we can use the weight vector $\bm{w}$ obtained from algorithm 2 to compute the $g$ values associated with each target data point as well as the $g$ value gap between different classes, which allows us to filter out classified data with low confidence. Experimental results for algorithm 2 are demonstrated in Appendix D.2
When applying optimal transport to solve domain adaptation problems, usually we do not have any label information for the target dataset nor the number of data points in each class. Therefore, the class imbalance between distributions could limit the performance of optimal transport based algorithms. In order for the OT map to transport target data clusters to correct source data clusters, we may need to rescale the weights of sampling points from the source data so that the total class weights are the same across source and target distributions. Under neural collapse type assumptions, we can provide theoretical guarantees that this issue can be solved automatically by selecting the "optimal weights". We provide the theory and numerical validation with synthetic data in Appendix C.1.1.
C.1.1 Unbalance Classes
@@
With the same notations of  8, suppose $\mu=p^{*}\mu_{1}+(1-p^{*})\mu_{2}$ for some $p^{*}\in(0,1)$. If $L_{i}\geq l_{i}+r_{\nu_{1}}+r_{\nu_{2}}+r_{\mu_{1}}+r_{\mu_{2}}$ then $\operatorname*{arg\,min}_{p\in[0,1]}W_{1}(\mu,\nu)=p^{*}$, where $\nu:=p\nu_{1}+(1-p)\nu_{2}$ for some $p\in(0,1)$.
@@
$\int_{F_{1}}\|T_{\nu}^{\mu}(x)-x\|d((\frac{1}{2}+\delta)\nu_{1})-b$
Since
$RHS\leq\delta r_{\nu_{1}}+\delta r_{\nu_{2}}\leq LHS,$
the optimality is proved. ∎
Figure 5: Unbalanced clusters with p=0.5
Figure 6: Unbalanced clusters
@@ Appendix D Experiment details
In our main experiments, we compare OT score with other confidence scores including Maxprob, Ent, Cossim, and JMDS. The details for other confidence scores are presented in Appendix A. We compare the performance of confidence scores on four standard UDA benchmarks: Digit recognition tasks, ImageCLEFDA, Office-Home, and VisDA-2017. All code can be efficiently executed on a single NVIDIA RTX 4070 GPU without requiring specialized hardware. For ImageCLEFDA, Office-Home, and VisDA-2017 datasets, we use ResNet-50 backbone pretrained on the ImageNet as a base network. Then we finetune the pretrained network on source domain for 100 epochs. For the finetuning, We use SGD optimizer with the momentum term set to be 0.9. We set lr=1e-4 for the base network and lr=1e-3 for the classifier layer. For digit recognition tasks, we use the VGG16 network as the base model. The network is initialized with random weights. We finetune this network on source domains using lr=1e-4, epochs=50, momentum=0.9, decay=1e-4. For OT score computation, we set the entropic regularization parameter $\varepsilon$ to be 0.5. A complete comparison for GMM pseudo labels is presented in Table 3.
Table 3: Evaluation of confidence scores based on AURC (GMM).
Dataset Task Maxprob Ent Cossim JMDS OT Score
Digits M $\rightarrow$ U 0.0055 0.0049 0.0051 0.0053 0.0033
U $\rightarrow$ M 0.0224 0.0228 0.0204 0.0355 0.0214
Avg. 0.0140 0.0139 0.0128 0.0204 0.0124
ImageCLEF-DA C $\rightarrow$ I 0.0319 0.0343 0.0386 0.0248 0.0221
C $\rightarrow$ P 0.2052 0.2078 0.1941 0.1395 0.1346
I $\rightarrow$ C 0.0088 0.0103 0.0076 0.0103 0.0079
I $\rightarrow$ P 0.1141 0.1131 0.1609 0.1227 0.1126
P $\rightarrow$ C 0.0094 0.0112 0.0079 0.0084 0.0090
P $\rightarrow$ I 0.0127 0.0141 0.0190 0.0131 0.0134
Avg. 0.0637 0.0651 0.0714 0.0531 0.0499
Office-Home Ar $\rightarrow$ Cl 0.2823 0.2860 0.2871 0.2825 0.2777
Ar $\rightarrow$ Pr 0.1320 0.1353 0.1219 0.1328 0.1313
Ar $\rightarrow$ Rw 0.0782 0.0835 0.1073 0.0820 0.0810
Cl $\rightarrow$ Ar 0.2421 0.2450 0.2503 0.2971 0.2218
Cl $\rightarrow$ Pr 0.1489 0.1517 0.1336 0.1708 0.1354
Cl $\rightarrow$ Rw 0.1366 0.1384 0.1412 0.1595 0.1241
Pr $\rightarrow$ Ar 0.2355 0.2453 0.2580 0.2715 0.2425
Pr $\rightarrow$ Cl 0.3417 0.3419 0.3180 0.3390 0.3257
Pr $\rightarrow$ Rw 0.0836 0.0875 0.1007 0.0952 0.0781
Rw $\rightarrow$ Ar 0.1439 0.1487 0.1785 0.1750 0.1391
Rw $\rightarrow$ Cl 0.2826 0.2812 0.2762 0.2760 0.2613
Rw $\rightarrow$ Pr 0.0693 0.0713 0.0725 0.0890 0.0637
Avg. 0.1814 0.1846 0.1871 0.1971 0.1735
VisDA-2017 T $\rightarrow$ V 0.3327 0.3312 0.3494 0.3090 0.2916
Pseudo-label generation via DSAN: To obtain pseudo labels, we need to further train the neural network using the DSAN algorithm with the following settings: number of training epochs = 20, transfer_loss_weight = 0.5, transfer_loss = LMMD, learning rate = 0.01, weight decay = $5\times 10^{-4}$, momentum = 0.9. lr_scheduler is enabled with lr_gamma = 0.0003, lr_decay = 0.75. A complete comparison for DSAN generated pseudo labels are provided in Table 4.
@@
D.1 Binary Toy Examples
To check the effectiveness of the OT score, we consider the binary classification DA task using USPS and MNIST datasets. When testing with the target dataset, we use both (1) the clean data and (2) the data set perturbed with Gaussian noise $\mathcal{N}(0,\sigma^{2}I)$, where we set $\sigma=0.4$ in our experiments. For each considered task, we compare (1) the accuracy using the pretrained NN on the source domain, (2) accuracy after using semi-discrete OT for domain adaptation, (3) accuracy after eliminating the $10\%$ data with lowest confidence estimated by the OT score. As shown as in Tables 5 and 6, SDOT-DA algorithm improves the performance of the classification network and the OT score can help us eliminate the uncertain data and further improve the accuracy. We also test on difficult pairs of digits and show the result in Table 7.
Tasks Methods $\sigma=0$ $\sigma=0.4$ Tasks Methods $\sigma=0$ $\sigma=0.4$
0, 1 Original 93.0 93.3 0, 6 Original 97.1 94.8
After OT 98.9 95.4 After OT 97.2 95.3
Quantile 0.1 99.9 96.8 Quantile 0.1 99.4 97.9
0, 2 Original 98.3 91.0 0, 7 Original 96.1 89.2
After OT 98.7 96.0 After OT 96.2 89.4
Quantile 0.1 100.0 98.8 Quantile 0.1 98.8 93.6
0, 3 Original 99.4 96.3 0, 8 Original 98.6 95.7
After OT 99.4 97.8 After OT 99.0 96.5
Quantile 0.1 100.0 99.7 Quantile 0.1 100.0 99.1
0, 4 Original 98.8 84.8 0, 9 Original 92.5 83.1
After OT 99.0 92.7 After OT 96.3 91.7
Quantile 0.1 100.0 96.0 Quantile 0.1 98.4 94.3
0, 5 Original 98.5 93.5 Average Original 96.9 91.3
After OT 98.8 96.3 After OT 98.2 94.6
Quantile 0.1 100.0 99.5 Quantile 0.1 99.6 97.3
Table 5: Accuracy for Binary Classification Tasks (USPS $\rightarrow$ MNIST)
Tasks Methods $\sigma=0$ $\sigma=0.4$ Tasks Methods $\sigma=0$ $\sigma=0.4$
0, 1 Original 99.8 92.4 0, 6 Original 99.4 92.8
After OT 99.8 95.5 After OT 99.6 93.8
Quantile 0.1 100.0 98.4 Quantile 0.1 100.0 97.7
0, 2 Original 99.4 94.2 0, 7 Original 70.9 70.9
After OT 99.4 96.5 After OT 99.0 97.2
Quantile 0.1 100.0 98.8 Quantile 0.1 99.8 99.1
0, 3 Original 99.2 96.5 0, 8 Original 98.6 83.6
After OT 99.6 97.7 After OT 99.2 92.6
Quantile 0.1 100.0 99.6 Quantile 0.1 100.0 96.5
0, 4 Original 99.1 93.9 0, 9 Original 99.6 92.7
After OT 99.8 94.3 After OT 99.6 94.0
Quantile 0.1 100.0 98.6 Quantile 0.1 100.0 98.5
0, 5 Original 99.2 89.9 Average Original 96.1 89.7
After OT 99.4 94.0 After OT 99.5 95.1
Quantile 0.1 100.0 97.9 Quantile 0.1 100.0 98.3
Table 6: Accuracy for Binary Classification Tasks (MNIST $\rightarrow$ USPS)
USPS $\rightarrow$ MNIST MNIST $\rightarrow$ USPS
Tasks Methods $\sigma=0$ $\sigma=0.4$ Tasks Methods $\sigma=0$ $\sigma=0.4$
0, 6 Original 97.1 94.8 0, 6 Original 99.4 92.8
After OT 97.2 95.3 After OT 99.6 93.8
Quantile 0.1 99.4 97.9 Quantile 0.1 100.0 97.7
0, 9 Original 92.5 83.1 0, 9 Original 99.6 92.7
After OT 96.3 91.7 After OT 99.6 94.0
Quantile 0.1 98.4 94.3 Quantile 0.1 100.0 98.5
1, 7 Original 87.7 79.7 1, 7 Original 64.2 64.2
After OT 96.6 85.0 After OT 98.7 91.5
Quantile 0.1 98.9 90.0 Quantile 0.1 100.0 95.6
2, 3 Original 98.7 94.2 2, 3 Original 99.4 94.5
After OT 98.8 95.2 After OT 99.7 95.6
Quantile 0.1 100.0 98.4 Quantile 0.1 100.0 98.8
2, 5 Original 98.5 95.7 2, 5 Original 99.4 93.0
After OT 98.8 96.2 After OT 99.7 93.8
Quantile 0.1 100.0 98.5 Quantile 0.1 100.0 98.4
3, 8 Original 98.4 88.5 3, 8 Original 99.0 93.7
After OT 98.8 93.9 After OT 99.6 93.7
Quantile 0.1 100.0 98.3 Quantile 0.1 100.0 98.3
4, 6 Original 98.9 89.2 4, 6 Original 98.9 93.2
After OT 99.0 92.8 After OT 99.1 95.1
Quantile 0.1 99.9 96.0 Quantile 0.1 99.6 97.3
5, 6 Original 98.3 84.6 5, 6 Original 99.4 93.3
After OT 98.5 92.5 After OT 99.4 93.6
Quantile 0.1 99.7 96.5 Quantile 0.1 100.0 97.5
6, 9 Original 96.3 83.4 6, 9 Original 99.7 89.6
After OT 98.3 94.7 After OT 100.0 96.2
Quantile 0.1 99.9 97.6 Quantile 0.1 100.0 99.7
Average Original 96.3 88.1 Average Original 95.4 89.7
After OT 98.0 93.0 After OT 99.5 94.1
Quantile 0.1 99.6 96.4 Quantile 0.1 99.9 98.0
Table 7: Accuracy for Hard Binary Classification Tasks
D.2 SDOT-DA
We test our SDOTDA algorithm on Digit recognition and Office-31 datasets. In our experiments, we assume the true proportions of each class. In the domain adaptation step, for each class in the source distribution, we use only the mean feature of that class for computation, which avoids the problem of storing the entire source dataset. As demonstrated from Tables 8 and 9, our SDOTDA algorithm significantly improves the accuracy on the given domain adaptation tasks.
Method $\mathcal{M}\rightarrow\mathcal{U}$ $\mathcal{U}\rightarrow\mathcal{M}$
GtA $92.8$ $90.8$
ADDA $89.4$ $90.1$
SRADA $94.1$ $\mathbf{98.0}$
RevGrad $77.1$ $73.0$
DRCN $91.8$ $73.7$
ETD $96.4$ $96.3$
Source Only $90.1$ $80.2$
SDOTDA $\mathbf{0.9711}$ $91.22$
Table 8: Accuracy for Classification Tasks Using Semi-Discrete OT(USPS $\rightarrow$ MNIST)
Method $\mathcal{A}\rightarrow\mathcal{W}$ $\mathcal{D}\rightarrow\mathcal{W}$ $\mathcal{W}\rightarrow\mathcal{D}$ $\mathcal{A}\rightarrow\mathcal{D}$ $\mathcal{D}\rightarrow\mathcal{A}$ $\mathcal{W}\rightarrow\mathcal{A}$ Average
Source Only $75.2$ $91.8$ $98.8$ $79.7$ $62.0$ $64.4$ 78.6
SDOTDA $82.6$ $95.0$ $99.4$ $87.8$ $65.9$ $66.6$ 82.9
Table 9: Office-31 Classification

### v2

@@ Appendix C Proofs
Now consider the optimal transport map $T_{\nu}^{\mu}$ restricted on $(E_{\nu_{1}}\cap F\cap B_{R})\cup(E_{\nu_{2}}\cap F)$. By (Villani et al., 2009, Theorem 4.6), this restricted map is an optimal transport map between the marginal measures.
@@
$\int_{A\cup B^{\prime}}\|\tilde{T}(x)-x\|\ dx\leq\delta(r_{\mu_{1}}+r_{\nu_{1}}+l_{1})+\delta(r_{\mu_{2}}+r_{\nu_{2}}+l_{2})<\delta(L_{1}+L_{2})\leq\int_{A\cup B^{\prime}}\|T_{\nu}^{\mu}(x)-x\|\ dx,$
@@
Choosing the difference CDC-D allows us to conclude $(T_{\bar{\nu}}^{\hat{\mu}})_{\#}\bar{\nu}_{1}=\hat{\mu}_{1}$ and $(T_{\bar{\nu}}^{\hat{\mu}})_{\#}\bar{\nu}_{2}=\hat{\mu}_{2}$ by setting $\bm{w}=[\bm{w}_{1},\bm{w}_{2}]$.
Conversely, let $\bm{w}$ be the corresponding weight vector of $T_{\bar{\nu}}^{\hat{\mu}}$ and assume $(T_{\bar{\nu}}^{\hat{\mu}})_{\#}\bar{\nu}_{1}=\hat{\mu}_{1}$, $(T_{\bar{\nu}}^{\hat{\mu}})_{\#}\bar{\nu}_{2}=\hat{\mu}_{2}$. Then $\bm{w}_{1}$ (or $\bm{w}_{2}$) differs from $\bm{m}$ (or $\bm{l}$) by some constant $C$ (or $D$) (Geiß et al., 2013, Theorem 2). By Theorem 6,
@@
This proposition shows how the classification accuracy improves with samples conditioned on high confidence scores $\Delta w$.
@@
Let $\nu_{1},\nu_{2}$ be the continuous probability measures with means $m_{1}$ and $m_{2}$, respectively and $\hat{\mu}_{i}$ consists of singletons $y_{i}$. Denote $\nu:=\frac{1}{2}\nu_{1}+\frac{1}{2}\nu_{2}$ and $\hat{\mu}:=\frac{1}{2}\hat{\mu}_{1}+\frac{1}{2}\hat{\mu}_{2}$. Suppose $\nu_{i}(|X_{i}-m_{i}|\geq t)\leq 2\exp\left(-\frac{t^{2}}{2\sigma^{2}}\right)$ and $\|m_{1}-y_{1}\|+\|m_{2}-y_{2}\|<\|m_{1}-y_{2}\|+\|m_{2}-y_{1}\|$, then $P\Big(T_{\nu}^{\hat{\mu}}(X_{i})\neq Y_{i}|g(X_{i})>\Delta w\Big)\leq 2\exp\left(-\frac{\min_{i=1,2}\operatorname{dist}(m_{i},\mathcal{S})^{2}}{2\sigma^{2}}\right)$, where
(1) $\mathcal{S}:=\Bigl\{x:\|x-y_{1}\|-(w^{*}+\Delta w)=\|x-y_{2}\|\Bigr\}$
(2) $d:=\|y_{2}-y_{1}\|,\qquad e:=\frac{y_{2}-y_{1}}{d},\qquad$ $m=\alpha e+u,\;u\perp e$ is the orthogal decomposition of $m$ and denote $\rho:=\|u\|$.
[…]
$\sqrt{t^{2}+r^{2}}\;=\;\sqrt{(t-d)^{2}+r^{2}}+(w^{*}+\Delta w),\qquad r\geq 0.$
@@
Let $w^{*}$ be the dual weight corresponding to $T_{\nu}^{\hat{\mu}}$ and let $w:=w+\Delta w$. Denote $\mathbb{L}_{\bm{w}}(y_{1}):=\Bigl\{x:\|x-y_{1}\|-w\leq\|x-y_{2}\|\Bigr\}$ and similarly for $\mathbb{L}_{\bm{w}}(y_{2})$.
Define $\mathcal{S}:=\Bigl\{x:\|x-y_{1}\|-w=\|x-y_{2}\|\Bigr\}$. Without loss of generality, we assume $y_{1}=\mathbf{0}$. For an arbitrary point $m\in\mathbb{R}^{n}$, write the orthogonal decomposition
$d:=\|y_{2}\|,\qquad e:=\frac{y_{2}}{d},\qquad m=\alpha e+u,\;u\perp e,\;\rho:=\|u\|.$
For every $x$ write
$x=t\,e+v,\qquad t\in\mathbb{R},\;v\perp e,\;r:=\|v\|.$
Under this decomposition
$\|x\|=\sqrt{t^{2}+r^{2}},\qquad\|x-y_{2}\|=\sqrt{(t-d)^{2}+r^{2}}.$
Hence $x\in\mathcal{S}$ iff
$\sqrt{t^{2}+r^{2}}\;=\;\sqrt{(t-d)^{2}+r^{2}}+w,\qquad r\geq 0.$ (8)
Since for any fixed $r$, $\sqrt{t^{2}+r^{2}}-\sqrt{(t-d)^{2}+r^{2}}$ is strictly increasing, solution to equation 8 is unique and we denote it by $t(r)$.
The squared distance between x=te+vx=t\,e+v and $m$ is
$\|x-m\|^{2}=(t-\alpha)^{2}+\|v-u\|^{2}=(t-\alpha)^{2}+r^{2}+\rho^{2}-2r\rho\cos\theta,$
where $\theta$ is the angle between $v$ and $u$. For fixed $(t,r)$ this expression is minimized when $\theta=0$, i.e. $v$ is chosen to be colinear with $u$. Without loss of generality set $v=(r/\rho)\,u$ when $\rho\neq 0$.
The minimal squared distance at any given $(t,r)$ is therefore $(t-\alpha)^{2}+(r-\rho)^{2}$. Since $t=t(r)$ is uniquely determined by $r$, the distance optimization reduces to
$\operatorname{dist}(m,\mathcal{S})=\min_{r\geq 0}\sqrt{\bigl(t(r)-\alpha\bigr)^{2}+\bigl(r-\rho\bigr)^{2}}.$
By a direct derivative analysis, the minimizer for $dist(m,\mathcal{S})$ is unique.
Therefore, take $m=m_{1}$, we have $\nu_{1}(\mathbb{L}_{\bm{w}}(y_{1}))\geq 1-2\exp\left(-\frac{\operatorname{dist}(m_{1},\mathcal{S})^{2}}{2\sigma^{2}}\right)$.
@@ Appendix D Experiment details
D.1 Synthetic Data
In this section, we use synthetic data to validate and illustrate our theoretical findings. Specifically, we consider a 2D scenario where data points are sampled from circular regions. The source domain consists of class-separated samples drawn from disjoint circles, whereas the target domain includes clusters with partial overlap. The distribution of the generated data is visualized in Figure LABEL:fig:overlap_samples. We compute the max-min values as described in Theorem 6 and present the results in Figure LABEL:fig:overlap_true. As shown in Figure LABEL:fig:overlap_predict, many of the generated pseudo labels within the overlapping region are misclassified. However, after removing low-confidence predictions, the remaining samples are almost entirely classified correctly, as illustrated in Figure LABEL:fig:overlap_pred_quantil_add. Notably, the separation between the two clusters becomes significantly more obvious after this filtering step.
(a) OT score performance on overlapping distributions.
D.2 Real-World Datasets
To ensure a fair comparison, we following the training setting of Lee et al. (2022). In our main experiments, we compare OT score with other confidence scores including Maxprob, Ent, Cossim, and JMDS. The details for other confidence scores are presented in Appendix B. We compare the performance of confidence scores on four standard UDA benchmarks: ImageCLEFDA, Office-Home, and VisDA-2017. All code can be efficiently executed on a single NVIDIA RTX 4070 GPU without requiring specialized hardware. For ImageCLEFDA, Office-Home datasets, we use ResNet-50 backbone pretrained on the ImageNet as a base network. The source model is trained for 50 epochs. For VisDA-2017, we use ResNet-101 for GMM pseudo labeling and ResNet-50 for DSAN pseudo labeling. The source model is obtained by finetuning a pretrained network on the source domain for 10 epochs. We use SGD optimizer with the momentum term set to be 0.9. We set lr=1e-4 for the base network and lr=1e-3 for the classifier layer. For digit recognition tasks, we use the ResNet-18 network as the base model. The network is initialized with random weights. We finetune this network on source domains using lr=1e-4, epochs=50, momentum=0.9, decay=1e-4. For OT score computation, we fix the entropic regularization parameter $\varepsilon$ to be 0.0001.
Pseudo-label generation via DSAN: To obtain pseudo labels, we need to further train the neural network using the DSAN algorithm with the following settings: number of training epochs = 20, transfer_loss_weight = 0.5, transfer_loss = LMMD, learning rate = 0.01, weight decay = $5\times 10^{-4}$, momentum = 0.9. lr_scheduler is enabled with lr_gamma = 0.0003, lr_decay = 0.75. comparison for DSAN generated pseudo labels are provided in Table 5.
We report mean ± standard deviation over three independent runs (random seeds) in Table 6 for Office-Home and Table 7 for VisDA-2017.
@@ D.2 Real-World Datasets
[…]
OTScore $58.0\pm 0.6$ $79.6\pm 0.1$ $81.5\pm 0.1$ $69.6\pm 0.4$ $80.2\pm 0.8$ $80.0\pm 0.2$
[…]
OTScore $68.3\pm 0.5$ $57.6\pm 0.7$ $82.3\pm 0.4$ $73.2\pm 0.1$ $61.1\pm 0.9$ $84.7\pm 0.4$ $73.0\pm 0.1$
Table 6: Accuracy (%) on Office-Home (ResNet-50).
Dataset Mean Accuracy Classwise Mean Accuracy
VisDA-2017 $85.0\pm 0.3$ $87.8\pm 0.1$
Table 7: Accuracy (%) on VisDA-2017.
Appendix E Unbalanced Classes
Theorem 11.
With the same notations of  9, suppose $\mu=p^{*}\mu_{1}+(1-p^{*})\mu_{2}$ for some $p^{*}\in(0,1)$. If $L_{i}\geq l_{i}+r_{\nu_{1}}+r_{\nu_{2}}+r_{\mu_{1}}+r_{\mu_{2}}$ then $\operatorname*{arg\,min}_{p\in[0,1]}W_{1}(\mu,\nu)=p^{*}$, where $\nu:=p\nu_{1}+(1-p)\nu_{2}$ for some $p\in(0,1)$.
Proof.
$\int_{F_{1}}\|T_{\nu}^{\mu}(x)-x\|d((\frac{1}{2}+\delta)\nu_{1})-b$
Since
$RHS\leq\delta r_{\nu_{1}}+\delta r_{\nu_{2}}\leq LHS,$
the optimality is proved. ∎
Figure 6: Unbalanced clusters with p=0.5
Figure 7: Unbalanced clusters